What are the dimensions of a rectangle with a perimeter of 44 units?
Answer by checkley77(12844)

THERE ARE MANY ANSWERS TO THIS QUESTION.
2L+2W=44
2(L+W)=44
L+W=44/2
L+W=22
ANY 2 POSITIVE INTEGERS WHOSE SUM=22 ARE SOLUTIONS.
11+11=22 ANS.
14+8=22 ANS.
3+19=22 ANS.
10+12=22 ANS.
5+17=22 ANS.